OpenStudy (anonymous):

NO ONE WILL ANSWER THIS QUESTION Let \[A _{1},A_{2},A_{3},....,A_{2n}\] be a finite AP with the sum of odd terms equals 24 and the sum of even terms equals to 30. If \[A_{2n}-A_{1}=10.5\] Find the number of terms.

OpenStudy (anonymous):

@Zekarias it must be 8

OpenStudy (anonymous):

right but how?

OpenStudy (anonymous):

actually last time i just gave u the value of n

OpenStudy (anonymous):

given A2n−A1=10.5 but A2n−A1=(2n-1)d hence 2nd -d =10.5 now (n/2)(A1 +A_n-1) =24 also (n/2)(A2+A_n) =30 subtracting these two we have (n/2) (A2 -A1 +A_n - A_n-1)=6 or (n/2)(d+d)= 6 or nd=6 but 2nd -d=10.5 or d=12-10.5 hence d=1.5 thus n=4 hence the number of terms being 2n =8

OpenStudy (anonymous):

I think its (n/2)(A1 +A_2n-1) =24 AND (n/2)(A2+A_2n) =30 ..... BUT THE ANSWER IS SAME. (A little typo error)

OpenStudy (anonymous):

yups and thanks